### v2

# Analytic Models for the Capacity Distribution in MDG-impaired Optical SDM Transmission

Abstract
In coupled space-division multiplexing (SDM) transmission systems, imperfections in optical amplifiers and passive devices introduce mode-dependent loss (MDL) and gain (MDG). These effects render the channel capacity stochastic and result in a decrease in average capacity. Several previous studies employ multi-section simulations to model the capacity of these systems. Additionally, relevant works derive analytically the capacity distribution for a single-mode system with polarization-dependent gain and loss (mode count $D=2$). However, to the best of our knowledge, analytic expressions of the capacity distribution for systems with $D>2$ have not been presented. In this paper, we provide analytic expressions for the capacity of SDM optical systems affected by MDG with arbitrary mode counts. The expressions rely on Gaussian approximations for the per-mode capacity distributions and for the overall capacity distribution, as well as on fitting parameters for the capacity cross-correlation among different modes. Compared to simulations, the derived analytical expressions exhibit a suitable level of accuracy across a wide range of practical scenarios.
Index Terms:
Space division multiplexing, channel capacity, channel models, optical fiber communications.
I Introduction
The increased capacity of multi-mode or multi-core fibers is a field of active research to handle the growing data traffic [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, unequal gains in amplifiers or unequal losses in passive devices and fibers generate an effect known as mode-dependent gain (MDG) and mode-dependent loss (MDL). In systems with MDG or MDL [11], the spatial coupling during fiber propagation turns the channel capacity into a random variable, introducing a probability of outage [12]. Communications systems with a stochastic signal-to-noise ratio (SNR) need to be designed to maintain the outage probability below accepted levels. To properly assess outages in the design of future space-division multiplexing (SDM) systems, statistical models for the capacity are necessary.
The capacity loss statistics for a set of amplification schemes are discussed in [13]. Expressions for the variance of the per-mode capacity loss for four different amplification schemes are derived. However, both per-mode capacity and total capacity distributions are not analyzed. Distributions for the per-mode modal gains have been presented in [14]. In addition, the paper obtains a closed-form expression for the average capacity. However, the capacity variance is not calculated. The impact of frequency diversity on the capacity distribution has been discussed in [15], but no analytical expression for the capacity distribution has been presented. The exact distribution for a single-mode-fiber optical system with polarization-dependent gain (PDG, $D=2$) has been presented in [16]. However, the statistics of SDM systems with higher mode counts have been addressed by simulation.
In this paper, we provide analytical expressions for the variance of the capacity distribution [^1] An analytical solution for the total capacity mean is presented in [14, Eq. (10)].. The derivations are based on the assumption of a Gaussian distribution for the per-mode and total capacity probability density functions (PDFs). The Gaussian assumption for the total capacity distribution has been discussed in [17], confirming its validity for increased mode counts. The joint capacity distribution is approximated as a multivariate normal, and the total capacity is obtained as the linear combination of the individuals. The derivations rely on the modal gain distributions outlined in [14].
II Instantaneous Capacity
Figure 1: (a) Logarithmic gain $\lambda_{\mathrm{dB}}$ distribution of each mode and ensemble, considering the sum of the linear gains equals to $D$. The ensemble distribution is normalized by $D$ in order to facilitate visual analysis. The per-mode distributions are obtained via simulation for ${D=6}$ and ${\sigma_{mdg}=5~{}\mathrm{dB}}$. (b) Instantaneous capacity distribution of each mode and total. The distributions are obtained via simulation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
Assuming that no channel state information (CSI) is available at the transmitter, the total instantaneous capacity of a narrowband SDM channel is given by [18]
$C_{T}=\mathrm{log}_{2}\left(\mathrm{det}\left[\mathbf{I}+\mathrm{SNR}\cdot\mathbf{H}\mathbf{H}^{H}\right]\right),$ (1)
where SNR is the ratio of total signal power and total noise power, $\mathbf{H}$ is the channel transfer matrix, and $\mathbf{I}$ is the identity matrix of size $D$, where $D$ is the mode count. The total instantaneous capacity can also be obtained from the eigenvalues $\lambda$ of the $\mathbf{H}\mathbf{H}^{H}$ [18] operator
$C_{T}=\sum_{i=1}^{D}C_{i}=\sum_{i=1}^{D}\mathrm{log}_{2}\left(1+\mathrm{SNR}\cdot\lambda_{i}\right),$ (2)
where $\lambda_{i}$ is the $i^{th}$ mode power gain experienced in the system. We consider noise as spatially white [^2] The spatial whiteness assumption has stronger validity for higher mode counts and wider bandwidths [19]., yielding equal noise power for all modes [14]. Assuming that SDM amplifiers maintain a constant total power after amplification, the sum of all $\lambda_{i}$ is set to $D$.
Equation (2) provides the instantaneous capacity for a single channel realization. In MDG-impaired systems, the random modal drift during fiber propagation results in a time-varying matrix $\mathbf{H}$, which in turn leads to fluctuations in both per-mode and total capacities. As an example, the modal-gain distribution for an optical transmission link affected by MDG is illustrated in Fig. 1(a). These gains result in a time-varying per-mode and total capacities that are illustrated in in Fig. 1(b). In [14], Ho et al. resort to random matrix theory to demonstrate that the overall PDF of an ensemble of all $\lambda_{i}$ eigenvalues, in decibels, is given by the well-known Gaussian unitary ensemble (GUE) spectral distribution, shown in Fig. 1(a) for $D=6$ [20, 21]. However, obtaining closed-form analytical models for the individual marginal $\lambda_{i}$ eigenvalue distributions is not trivial, except for specific cases, such as low eigenvalue counts (${D\leq 3}$[22], or for the smallest and largest eigenvalues as specified by the Tracy-Widom distribution [23].
III Per-Mode Individual Capacity Distributions
We evaluate the statistics of the total capacity $C_{T}$ based on the per-mode individual $\lambda_{i}$ and $C_{i}$ distributions. Approximations on $C_{i}$ distributions are obtained from two methods derived from the GUE spectral and Wigner semicircular distribution for $\lambda_{i}$.
III-A Approximations on the $C_{i}$ distribution based on the GUE spectral distribution for $\lambda_{i}$
The distribution of the $\lambda_{\mathrm{dB}}$ eigenvalues for a $D^{th}$ order GUE is given by [14, Table 1], [22, Eq. (S.15)]
$\begin{split}f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})=&\frac{\alpha_{\lambda_{\mathrm{dB}},D}}{\sigma_{\mathrm{mdg}}}e^{-\frac{(D+1)}{2}\left(\frac{\lambda_{\mathrm{dB}}-\mu_{\lambda_{\mathrm{dB}}}}{\sigma_{\mathrm{mdg}}}\right)^{2}}\\ &\sum_{k=0}^{D-1}\beta_{\lambda_{\mathrm{dB}},D,k}\left(\frac{\lambda_{\mathrm{dB}}-\mu_{\lambda_{\mathrm{dB}}}}{\sigma_{\mathrm{mdg}}}\right)^{2k}\end{split}$ (3)
where $\alpha_{\lambda_{\mathrm{dB}},D}$ is a normalization factor such that the PDF integral equals 1, and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are polynomial coefficients. $\alpha_{\lambda_{\mathrm{dB}},D}$ and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are constant for a given number of modes. For $D\leq 8$ the values of $\alpha_{\lambda_{\mathrm{dB}},D}$ and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are provided in [14, Table 1], and the derivation of the coefficients is presented in appendix A. Values [^3] The notation $\sigma_{\mathrm{mdg}}$ was preferred to $\sigma_{\lambda_{\mathrm{dB}}}$, as the former is predominant in literature. $\sigma_{\mathrm{mdg}}$ and $\mu_{\lambda_{\mathrm{dB}}}$ are the standard deviation and the mean of $f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})$, respectively. Value $\mu_{\lambda_{\mathrm{dB}}}$ is given such that the linearised eigenvalue distribution $f_{\lambda}(\lambda)$ has unitary mean, and is given by (see appendix B)
[…]
where $\widehat{f}_{\lambda_{\mathrm{dB}}}(x)$ is the zero mean $\lambda_{\mathrm{dB}}$ distribution.
We assume that the means of individual eigenvalues $\lambda_{\mathrm{dB},i}$ are near the local maxima of $f_{\lambda_{\mathrm{dB}}}$. We further assume that $\lambda_{\mathrm{dB},i}$ is Gaussian-distributed with standard deviation $\sigma_{\lambda_{dB,i}}$ and mean $\mu_{\lambda_{\mathrm{dB},i}}$. The ordering of the indices is given by their mean value such that $\mu_{\lambda_{\mathrm{dB},1}}<\mu_{\lambda_{\mathrm{dB},2}}<\cdots<\mu_{\lambda_{\mathrm{dB},D}}$.
We approximate $\mu_{\lambda_{\mathrm{dB},i}}$ by the odd-index zeros of the $f_{\lambda_{\mathrm{dB}}}$ derivative (for a detailed derivation, see appendix C), yielding
$\begin{split}\sum_{k=0}^{D-1}\frac{\beta_{\lambda_{\mathrm{dB}},D,k}}{\sigma_{\mathrm{mdg}}^{2k}}\left[2k\cdot\left(\mu_{\lambda_{\mathrm{dB},i}}-\mu_{\lambda_{\mathrm{dB}}}\right)^{2k-1}\right.\\ \left.-\frac{D+1}{\sigma_{\mathrm{mdg}}^{2}}\left(\mu_{\lambda_{\mathrm{dB},i}}-\mu_{\lambda_{\mathrm{dB}}}\right)^{2k+1}\right]&=0.\end{split}$ (5)
Figure 2: Analytical approximation of per-mode capacity mean (a) and standard deviation (b) for ${\mathrm{SNR}=10~{}\mathrm{dB}}$ and ${\sigma_{mdg}=5~{}\mathrm{dB}}$. For ${D\leq 8}$ the GUE distribution method is considered. Higher mode counts use the Wigner semicircular CDF method.
Figure 3: Gaussian per-mode capacity PDF s for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
The values of $\sigma_{\lambda_{\mathrm{dB},i}}$ are obtained by making the ensemble and individual PDF values equal at $\mu_{\lambda_{\mathrm{dB},i}}$ and applying a normalization factor of $1/D$
$\sigma_{\lambda_{\mathrm{dB},i}}=\frac{1}{D\sqrt{2\pi}f_{\lambda_{\mathrm{dB}}}(\mu_{\lambda_{\mathrm{dB},i}})}.$ (6)
From the obtained $\mu_{\lambda_{\mathrm{dB},i}}$ and $\sigma_{\lambda_{\mathrm{dB},i}}$, the per-mode PDF is approximated by a Gaussian distribution
$f_{\lambda_{\mathrm{dB},i}}(\lambda_{\mathrm{dB},i})=\frac{1}{\sigma_{\lambda_{\mathrm{dB},i}}\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{\lambda_{\mathrm{dB},i}-\mu_{\lambda_{\mathrm{dB},i}}}{\sigma_{\lambda_{\mathrm{dB},i}}}\right)^{2}}.$ (7)
The relation between the individual eigenvalue distribution $f_{\lambda_{\mathrm{dB},i}}$ and its corresponding capacity distribution $f_{C_{i}}(c_{i})$ is obtained by transformation as
$f_{C_{i}}(c_{i})=\frac{10\cdot\mathrm{ln}\left(2\right)2^{c_{i}}}{\mathrm{ln}\left(10\right)\left(2^{c_{i}}-1\right)}f_{\lambda_{\mathrm{dB},i}}\left(\frac{10}{\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{c_{i}}-1}{\mathrm{SNR}}\right)\right).$ (8)
We also approximate the $C_{i}$ PDF by a Gaussian distribution with standard deviation $\sigma_{C_{i}}$ and mean $\mu_{C_{i}}$. Setting the derivative of (8) to zero (see appendix D), the value of $\mu_{C_{i}}$ can be obtained as the solution of the simplified equation
$\frac{10\cdot 2^{\mu_{C_{i}}}}{\sigma_{\lambda_{\mathrm{dB},i}}^{2}\mathrm{ln}\left(10\right)}\left(\frac{10}{\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{\mu_{C_{i}}}-1}{\mathrm{SNR}}\right)-\mu_{\lambda_{\mathrm{dB},i}}\right)+1=0.$ (9)
The standard deviation values $\sigma_{C,i}$ are obtained by assuming a Gaussian approximation and setting (8) equal at $\mu_{C,i}$, which is achieved for
$\sigma_{C_{i}}=\frac{1}{\sqrt{2\pi}f_{C_{i}}(\mu_{C_{i}})}.$ (10)
An evaluation of $\mu_{C_{i}}$ and $\sigma_{C_{i}}$ for $2\leq D\leq 8$ is presented in Fig. 2, exhibiting an excellent agreement. The accuracy of the Gaussian approximation for the individual capacity distributions $f_{C_{i}}$, for a system with $D=6$, $\mathrm{SNR}=10~{}\mathrm{dB}$ and $\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}$, is shown in Fig. 3. The results reveal, again, a suitable agreement between simulation and the derived analytical curves.
III-B Approximations on the $C_{i}$ distribution based on the Wigner semicircular distribution for $\lambda_{i}$
The limiting distribution of $\lambda_{\mathrm{dB}}$ as $D\rightarrow\infty$ is given by the Wigner semicircular distribution as [20, I][14, Table 1]
$\begin{gathered}f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})=\frac{1}{2\pi\sigma_{\mathrm{mdg}}}\sqrt{4-\frac{\left(\lambda_{\mathrm{dB}}-\mu_{\lambda_{\mathrm{dB}}}\right)^{2}}{\sigma_{\mathrm{mdg}}^{2}}},\\ -2\sigma_{\mathrm{mdg}}+\mu_{\lambda_{\mathrm{dB}}}\leq\lambda_{\mathrm{dB}}\leq 2\sigma_{\mathrm{mdg}}+\mu_{\lambda_{\mathrm{dB}}},\end{gathered}$ (11)
with $\mu_{\lambda_{\mathrm{dB}}}$ given in (4).
For this limiting case, the capacity PDF and CDF for the ensemble of eigenvalues can be derived analytically. The capacity PDF is obtained applying (11) into (8) and is given by
$\begin{split}f_{C}(c)=&\frac{5\cdot\mathrm{ln}\left(2\right)2^{c}}{\pi\sigma_{\mathrm{mdg}}\mathrm{ln}\left(10\right)\left(2^{c}-1\right)}\\ &\sqrt{4-\frac{1}{\sigma^{2}_{\mathrm{mdg}}}\left(\frac{10}{\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{c}-1}{\mathrm{SNR}}\right)-\mu_{\lambda_{\mathrm{dB}}}\right)^{2}},\end{split}$ (12)
and the CDF, developed in further detail in appendix E, is given by
$\begin{split}F_{C}(c)=&\frac{1}{2}+\left(\frac{5}{\pi\sigma_{\mathrm{mdg}}\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{c}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{2\pi\sigma_{\mathrm{mdg}}}\right)\\ &\sqrt{1-\left(\frac{5}{\sigma_{\mathrm{mdg}}\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{c}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{2\sigma_{\mathrm{mdg}}}\right)^{2}}\\ &+\frac{1}{\pi}\mathrm{sin^{-1}}\left(\frac{5}{\sigma_{\mathrm{mdg}}\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{c}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{2\sigma_{\mathrm{mdg}}}\right),\end{split}$ (13)
where both distributions are limited in the following support
$\begin{split}\mathrm{log}_{2}\left(\mathrm{SNR}\cdot 10^{\frac{\mu_{\lambda_{\mathrm{dB}}}-2\sigma_{\mathrm{mdg}}}{10}}+1\right)\leq c\\ \leq\mathrm{log}_{2}\left(\mathrm{SNR}\cdot 10^{\frac{\mu_{\lambda_{\mathrm{dB}}}+2\sigma_{\mathrm{mdg}}}{10}}+1\right).\end{split}$ (14)
Figure 4: Capacity CDF and ${\mu_{C_{i}}}$ estimation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$.
Assuming that the individual PDF s are symmetric and have minimal overlap, the values of $\mu_{C_{i}}$ can be obtained from the CDF of the ensemble as
$\mu_{C_{i}}=F^{-1}_{C}\left(\frac{i-1/2}{D}\right),$ (15)
where $F^{-1}_{C}$ is the inverse function of the capacity CDF.
An assessment of the CDF-derived approximation can be found in Fig. 4, for a system with $D=6$, $\mathrm{SNR}=10~{}\mathrm{dB}$ and $\sigma_{mdg}=5~{}\mathrm{dB}$, showing excellent agreement.
The values of $\sigma_{C_{i}}$ can then be obtained from the ensemble PDF and $\mu_{C_{i}}$, such as in (6), yielding
$\sigma_{C_{i}}=\frac{1}{D\sqrt{2\pi}f_{C}(\mu_{C_{i}})}.$ (16)
Differences between the predictions generated by the Wigner semicircular distribution and the GUE spectral are expected. The Wigner semicircular distribution flattens the local maxima of the GUE spectral distribution. However, the divergences fade as the mode count increases. In Fig. 2, the predictions for $D=12$, $D=20$ and $D=40$ are obtained based on the Wigner semicircle distribution. A significant $\sigma_{C_{i}}$ error is observed at $D=12$, attributed to the divergence of the Wigner semicircular distribution from the actual distribution. This error diminishes as the number of modes increases.
IV Total Capacity Distribution
Owing to correlations between $\lambda_{i}$ variables, the individual distributions are not sufficient to model the statistical parameters of the total capacity. As seen in Fig. 3, the distributions of per-mode individual capacities closely approximate normal distributions. Therefore, in this work, we also propose to approximate the total capacity PDF by a multivariate normal distribution. A sum of individual elements of a multivariate normal distribution has a normal distribution given by
$f_{C_{T}}(c_{T})=\frac{1}{\sigma_{C_{T}}\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{c_{T}-\mu_{C_{T}}}{\sigma_{C_{T}}}\right)^{2}},$ (17)
where the variance $\sigma_{C_{T}}^{2}$ and mean $\mu_{C_{T}}$ can be obtained, respectively, by
$\begin{split}\sigma^{2}_{C_{T}}&=\sum_{i,j=1,1}^{D,D}\sigma_{C_{i}}\sigma_{C_{j}}\rho_{i,j}\\ \mu_{C_{T}}&=\sum_{i=1}^{D}\mu_{C_{i}}.\end{split}$ (18)
Alternatively, an already proposed exact solution for $\mu_{C_{T}}$ has been proposed in [14, Eq. (10)][16, Eq. (12)]
[…]
where $\widetilde{f}_{\lambda_{\mathrm{dB}}}(x)$ is the unitary variance, zero mean $\lambda_{\mathrm{dB}}$ distribution.
In (18), $\rho_{i,j}$ is the correlation coefficient between the $i^{th}$ and $j^{th}$ capacity PDF s. The correlation coefficient is related to the covariance by
$\rho_{i,j}=\frac{\mathrm{cov}\left(C_{i},C_{j}\right)}{\sigma_{C_{i}}\sigma_{C_{j}}}.$ (20)
Figure 5: Correlation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
TABLE I: Correlation function coefficients. The results were obtained with $10^{-10}$ $\gamma_{0}$ error tolerance and $10^{4}$ $\gamma_{1}$ iterations.
$\mathrm{SNR}=5\ \mathrm{dB}$ $\mathrm{SNR}=10\ \mathrm{dB}$ $\mathrm{SNR}=15\ \mathrm{dB}$ $\mathrm{SNR}=20\ \mathrm{dB}$
$D$ $\gamma_{0}$ $\gamma_{1}$ $\gamma_{0}$ $\gamma_{1}$ $\gamma_{0}$ $\gamma_{1}$ $\gamma_{0}$ $\gamma_{1}$
Method III-A 2 $2.16297989$ $-1.424485\mathrm{E-}3$ $2.16381222$ $-3.241463\mathrm{E-}4$ $2.16393724$ $-5.664255\mathrm{E-}5$ $2.16395171$ $-1.031541\mathrm{E-}5$
3 $1.09160893$ $4~{}~{}6.095528\mathrm{E-}$ 4 $1.09153003$ $4~{}~{}1.498233\mathrm{E-}$ 4 $1.09149336$ $5~{}~{}4.022044\mathrm{E-}$ 5 $1.09147997$ $6~{}~{}7.691563\mathrm{E-}$ 6
4 $0.72791071$ $4~{}~{}1.114647\mathrm{E-}$ 4 $0.72795167$ $5~{}~{}5.751291\mathrm{E-}$ 5 $0.72793629$ $5~{}~{}2.789583\mathrm{E-}$ 5 $0.72792665$ $6~{}~{}6.419168\mathrm{E-}$ 6
5 $0.54480425$ $-1.547994\mathrm{E-}4$ $0.54492813$ $-2.174217\mathrm{E-}5$ $0.54493125$ $5~{}~{}1.038573\mathrm{E-}$ 5 $0.54492585$ $6~{}~{}5.543675\mathrm{E-}$ 6
6 $0.43511438$ $5~{}~{}7.630925\mathrm{E-}$ 5 $0.43513127$ $5~{}~{}3.758373\mathrm{E-}$ 5 $0.43512306$ $5~{}~{}1.614824\mathrm{E-}$ 5 $0.43511820$ $6~{}~{}3.966745\mathrm{E-}$ 6
7 $0.36203898$ $-1.120187\mathrm{E-}4$ $0.36211666$ $-1.754819\mathrm{E-}5$ $0.36212190$ $6~{}~{}4.308864\mathrm{E-}$ 6 $0.36212014$ $6~{}~{}3.102686\mathrm{E-}$ 6
8 $0.31008496$ $-1.768119\mathrm{E-}4$ $0.31016253$ $-3.614410\mathrm{E-}5$ $0.31017045$ $-2.168861\mathrm{E-}6$ $0.31017007$ $6~{}~{}1.195690\mathrm{E-}$ 6
Method III-B 9 $0.27114473$ $4~{}~{}1.871205\mathrm{E-}$ 4 $0.27107389$ $5~{}~{}3.189872\mathrm{E-}$ 5 $0.27106495$ $6~{}~{}1.572890\mathrm{E-}$ 6 $0.27106446$ $-4.804128\mathrm{E-}7$
10 $0.24115580$ $5~{}~{}8.250165\mathrm{E-}$ 5 $0.24113560$ $6~{}~{}3.632280\mathrm{E-}$ 6 $0.24113577$ $-4.177426\mathrm{E-}6$ $0.24113686$ $-1.934882\mathrm{E-}6$
12 $0.19777600$ $5~{}~{}7.311636\mathrm{E-}$ 5 $0.19775961$ $-1.298120\mathrm{E-}7$ $0.19776223$ $-7.621126\mathrm{E-}6$ $0.19776444$ $-3.509853\mathrm{E-}6$
20 $0.11577972$ $-2.081275\mathrm{E-}4$ $0.11585818$ $-7.543138\mathrm{E-}5$ $0.11587780$ $-2.177745\mathrm{E-}5$ $0.11588300$ $-6.367806\mathrm{E-}6$
30 $0.07681464$ $-2.282790\mathrm{E-}4$ $0.07689957$ $-8.327005\mathrm{E-}5$ $0.07691975$ $-2.324244\mathrm{E-}5$ $0.07692502$ $-6.536148\mathrm{E-}6$
40 $0.05768991$ $-1.842371\mathrm{E-}4$ $0.05776636$ $-6.941402\mathrm{E-}5$ $0.05778465$ $-2.036533\mathrm{E-}5$ $0.05778933$ $-5.994924\mathrm{E-}6$
Figure 6: Estimation of the total capacity variance with the analytical model compared to simulated values. For $4$ (a) and $8$ modes (b), the analytical method relies on the individual statistical values derived from the GUE spectral distribution, as discussed in Section III-A. For $12$ (c) and $40$ (d) modes, values are derived from the Wigner semicircular distribution, as discussed in Section III-B.
Figure 7: Approximation error for ${\mathrm{SNR}=10~{}\mathrm{dB}}$. For comparison fairness, all the per-mode individual values are obtained from the Wigner semicircular CDF method.
An exact equation for $\rho_{i,j}$ is unknown. However, as observed in Fig. 5, a suitable empirical approximation is given by
$\rho_{i,j}=e^{-|i-j|}+\left(e^{-|i-j|}-1\right)\left(\gamma_{0}+\gamma_{1}\cdot\sigma_{mdg}^{2.75}\right),$ (21)
where the coefficients $\gamma_{0}$ and $\gamma_{1}$ are specific to a given mode count and SNR value. In this paper, we numerically obtain the values of $\gamma_{0}$ and $\gamma_{1}$ such that the mean squared logarithmic error (MSLE) between analytical and simulation capacity variance is minimized, i.e.,
$\left(\gamma_{0},\gamma_{1}\right)=\mathrm{argmin}_{\gamma_{0},\gamma_{1}}\left(\mathbb{E}\left\{\left[\mathrm{ln}\left(\widehat{\sigma}^{2}_{C_{T}}\right)-\mathrm{ln}\left(\sigma^{2}_{C_{T}}\right)\right]^{2}\right\}\right).$ (22)
The methodology to obtain the coefficients is performed iteratively until a tolerance threshold is reached. The algorithm implemented in this work firstly updates $\gamma_{0}$ until the error for the lowest considered $\sigma_{\mathrm{mdg}}$ is smaller than a defined tolerance. The value of $\gamma_{1}$ is updated considering the MSLE for the entire evaluated $\sigma_{\mathrm{mdg}}$ range, with small corrections to $\gamma_{0}$ if the resulting approximation function is not monotonic increasing. We select the logarithmic error because the capacity variance scales exponentially with $\sigma_{mdg}$ when measured in decibels. The coefficients for a practical range of mode counts and SNR are given in Table I.
An evaluation of the analytical method is presented in Fig. 6. The results indicate an accurate variance approximation for SNR values lower than $20~{}\mathrm{dB}$ for all evaluated mode counts. For $\textrm{SNR}=20~{}\mathrm{dB}$, there is a higher divergence, particularly with the Wigner semicircular CDF method. The source of estimation discrepancy likely stems from the per-mode individual distributions estimated from the capacity CDF, as well as approximations made in the correlation function. Fig. 7 evaluates the accuracy of the method based on Wigner semicircular CDF for different values of $\sigma_{mdg}$ and mode counts. The results evidence accurate estimates for mode counts higher than 20. Fig. 8 compares the simulated total capacity distribution with the analytical distributions produced with the Wigner semicircular CDF method. The presented Gaussian approximation exhibits a good agreement with the simulated curve.
Figure 8: Total Gaussian PDF for ${D=20}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
V Impact of frequency diversity
In MDG-impaired systems, the random coupling among modes turns the capacity into a random variable, eventually generating outages. As demonstrated in [15], modal dispersion generates a frequency-dependent channel frequency response that mitigates the impact of deep fades. This effect, known as frequency diversity, benefits the outage capacity [24]. As seen in Fig. 9, the total capacity standard deviation has a weak dependence on the mode count, in agreement with [24]. This effect has been called mode diversity in [16]. According to the effect of mode diversity, the outage probability is more significant in lower mode counts, as fewer paths are available to mitigate outages.
Figure 9: Total capacity deviation per mode count for ${\sigma_{mdg}=5~{}\mathrm{dB}}$ from the analytical equations.
So far, this paper considered a narrowband channel with a flat frequency response. However, as presented in [15], modal dispersion induces frequency-dependent mode gains. In this case, the total capacity in a given bandwidth can be calculated by dividing the spectrum into $N$ narrowband frequency bins and adding the corresponding capacities [16]. In this model, the channel response of a given frequency bin is independent of that in adjacent bins [15]. Considering the distributions presented in Section III, the averaged capacity variance of the $i^{th}$ mode, considering $N$ independent frequency bins is given by
$\overline{\sigma}^{2}_{C_{i}}=\frac{\sigma^{2}_{C_{i}}}{N}.$ (23)
Substituting the individual capacity variances in (18) with (23), we see that the average total capacity variance is reduced by the same rate of $1/N$.
VI Case Study
Assuming strong-mode-coupling, the formulas presented in this paper can be used for system design if the accumulated MDG and optical SNR are known to the operator. Estimation techniques for both parameters are discussed in [25] and [26]. As a case study, we calculate the outage capacity of a 3-mode polarization multiplexed link ($D=6$) with $\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}$ and $\mathrm{SNR}=10~{}\mathrm{dB}$, for an outage probability $p_{\mathrm{out}}=1\%$. We use the GUE spectral distribution method presented in Section III-A.
The mean logarithmic power gain, considering an amplification scheme with total power control, is obtained from (4) by numeric integration as
$\mu_{\lambda_{\mathrm{dB}}}=-2.609~{}\mathrm{dB}.$ (24)
The $\lambda_{\mathrm{dB}}$ distribution coefficients can be obtained from [14, Table 1] or derived analytically, as detailed in appendix A. They are given by
$\begin{split}\alpha_{\lambda_{\mathrm{dB}},6}=&\sqrt{\frac{14}{\pi}}\\ \boldsymbol{\beta}_{\lambda_{\mathrm{dB}},6}=&\left[\frac{322}{3125},~{}\frac{4557}{1250},~{}\frac{-17493}{625},~{}\right.\\ &\left.~{}\frac{256221}{3125},~{}\frac{-259308}{3125},~{}\frac{453789}{15625}\right].\end{split}$ (25)
The individual capacities mean and standard deviation are calculated from (5) and (6), respectively, and are given by
$\begin{split}\boldsymbol{\mu}_{C}=&[1.022,~{}1.653,~{}2.330,~{}3.067,~{}3.887,~{}4.865],~{}[b/s/Hz]\\ \boldsymbol{\sigma}_{C}=&[0.192,~{}0.202,~{}0.217,~{}0.238,~{}0.276,~{}0.362],~{}[b/s/Hz],\end{split}$ (26)
where the values of $\boldsymbol{\mu}_{C}$ are obtained numerically, from the odd-index crossings of (5) along the $x$ axis.
The individual capacity distribution for the given parameters is presented in Fig. 3.
The correlation coefficients are obtained from (21), considering $\gamma_{0}=0.43513127$ and $\gamma_{1}=3.758373\cdot 10^{-5}$, as provided in table I, and are given in matrix form as
$\boldsymbol{\rho}=\begin{bmatrix}\hskip 7.11317pt1&0.091&-0.244&-0.367&-0.412&-0.429\\ \hskip 7.11317pt0.091&1&0.091&-0.244&-0.367&-0.412\\ -0.244&0.091&1&0.091&-0.244&-0.367\\ -0.367&-0.244&0.091&1&0.091&-0.244\\ -0.412&-0.367&-0.244&0.091&1&0.091\\ -0.429&-0.412&-0.367&-0.244&0.091&1\\ \end{bmatrix}$ (27)
The total capacity statistical parameters are obtained from (18) and are given by
$\begin{split}\mu_{C_{T}}=\hskip 2.84526pt&16.825,~{}[b/s/Hz]\\ \sigma_{C_{T}}=\hskip 2.84526pt&0.181,~{}[b/s/Hz].\end{split}$ (28)
The total capacity distribution for the given parameters can be seen in Fig. 10.
Considering the Gaussian model, the outage capacity $c_{\mathrm{out}}$ can be obtained from the Gaussian CDF and the outage probability as
$c_{\mathrm{out}}=F_{C_{T}}^{-1}(p_{\mathrm{out}})=\sqrt{2}\sigma_{C_{T}}\mathrm{erf}^{-1}\left(2p_{\mathrm{out}}-1\right)+\mu_{C_{T}},$ (29)
where $\mathrm{erf}^{-1}(\cdot)$ is the inverse error function.
For the defined parameters the outage capacity is
$c_{\mathrm{out}}=16.527,~{}[b/s/Hz].$ (30)
Considering the total capacity mean presented in (18), the outage capacity is marginally underestimated. Using (19), the outage capacity is more accurately estimated as
$c_{\mathrm{out},\mathrm{int}}=16.654,~{}[b/s/Hz].$ (31)
In the presence of frequency diversity of $N=2$, the average total capacity deviation is reduced and is given by
$\overline{\sigma}_{C_{T}}=0.128,~{}[b/s/Hz].$ (32)
Considering the mean capacity given by (18), the reduced variance increases the average outage capacity to
$\overline{c}_{\mathrm{out}}=16.614,~{}[b/s/Hz].$ (33)
Figure 10: Total Gaussian PDF for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$. For the frequency diverse case, we consider 2 uncorrelated frequency bins. The estimated outage capacities consider an outage probability of ${p_{\mathrm{out}}=1\%}$.
VII Conclusion
We propose statistical models that approximate the individual per-mode and total capacity variances for strongly-coupled SDM transmission systems. The model is based on the Gaussian assumption for the individual per-mode capacity. The total capacity distribution is modeled as a joint multivariate normal distribution given by a linear combination of the individuals. The individual per-mode capacity distributions are derived using two methods: one based on the GUE spectral distribution for channel gains, and the other based on the Wigner semicircular distribution. The correlation between per-mode individual capacities is addressed by an empirical analytical model with fitting coefficients obtained by numerical simulations. The analytical results exhibit a suitable agreement with simulations in the investigated scenarios. The narrowband results are extended to arbitrary bandwidths resorting to existing theory on frequency diversity.
Appendix A Distribution of the ensemble of $\lambda_{\mathrm{dB}}$ eigenvalues
As described in [22], the logarithm power gain $\lambda_{\mathrm{dB}}$ follows the fixed-trace GUE. The joint PDF of the GUE spectral distribution is given by [27, Eq. (3.3.7)]
$f_{\boldsymbol{\lambda}}(\boldsymbol{\lambda})=\alpha_{\boldsymbol{\lambda},D}\exp\left(-\sum_{k=1}^{D}\lambda_{k}^{2}\right)\prod_{\begin{subarray}{c}j,i\in[1,D]\\ i<j\end{subarray}}(\lambda_{i}-\lambda_{j})^{2},$ (34)
with $\lambda_{1}<\lambda_{2}<\cdots<\lambda_{D}$, and where $\alpha_{\boldsymbol{\lambda},D}$ is a normalization factor, such that the PDF integral is unitary.
The fixed-trace GUE has the additional constraint that the sum of the eigenvalues is constant. As pointed out by [28], the fixed-trace is a property of stable systems. Considering a GUE matrix $\mathbf{A}\in\mathbb{C}^{D\times D}$, the fixed-trace condition specifies that
$\text{tr}(\mathbf{A})=\sum_{i=1}^{D}\lambda_{i}=\kappa,$ (35)
where $\text{tr}(\cdot)$ represents the matrix trace and $\kappa$ is a constant.
The ensemble spectral distribution, related to any eigenvalue $\lambda$ in set $\lambda_{i},i\in[1,D]$, is obtained from the $k$-point correlation function in the case of $k=1$, defined in [29, III] and [27, Eq. (6.1.2)].
The fixed-trace solution of the ensemble spectral distribution normalized to zero mean is developed in detail in [22, From (S.3) to (S.15)]. Defining an auxiliary variable $t$ where its $n^{th}$ order polynomial is expressed in terms of $\lambda$ as
$t^{n}\xmapsto{}\left[(-1)^{n}H_{n}\left(\lambda\sqrt{\frac{D(D+1)}{2}}\right)\right],$ (36)
the fixed-trace solution normalized to unitary variance is given by
$f_{\lambda}(\lambda)=\frac{\sqrt{D+1}}{D\sqrt{2\pi}}e^{-\frac{(D+1)}{2}\lambda^{2}}\sum^{D-1}_{k=0}\frac{1}{2^{k}k!}H^{2}_{k}\left(\frac{t}{2\sqrt{D-1}}\right),$ (37)
where $H_{n}(\cdot)$ is the physicist’s Hermite polynomial given by
$H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}=n!\sum_{k=0}^{\lfloor n/2\rceil}(-1)^{k}\frac{(2x)^{n-2k}}{k!(n-2k)!},$ (38)
[…]
Expanding the Hermite polynomial in (38) and replacing the auxiliary variable $t$ as defined in (37), the logarithmic gain distribution, with proper mean and standard deviation scaling, can be rewritten as (3).
As pointed out by [20, I], at the limit $D\rightarrow\infty$ the distribution given in (38) converges to
$\lim_{D\rightarrow\infty}f_{\lambda}(\lambda)=\frac{1}{2\pi}\sqrt{4-\lambda^{2}},~{}2\leq\lambda\leq 2,$ (39)
and with proper mean and standard deviation scaling, is rewritten as (11).
Appendix B Equation for the mean $\mu_{\lambda_{\mathrm{dB}}}$ of the linearized eigenvalues
Given $\lambda_{\mathrm{dB}}=10\cdot\mathrm{log}_{10}\left(\lambda\right)$, the linearized eigenvalue distribution is related to the $\lambda_{\mathrm{dB}}$ distribution by
$f_{\lambda}(\lambda)=\frac{10}{\mathrm{ln}\left(10\right)\lambda}f_{\lambda_{\mathrm{dB}}}\left(10\cdot\mathrm{log}_{10}\left(\lambda\right)\right),$ (40)
and the $\lambda$ mean is given by
[…]
where $\widehat{f}_{\lambda_{\mathrm{dB}}}(\cdot)$ is the $\lambda_{\mathrm{dB}}$ distribution shifted by $\mu_{\lambda_{\mathrm{dB}}}$, resulting in a zero mean PDF. Using $x=10\cdot\mathrm{log}_{10}\left(\lambda\right)-\mu_{\lambda_{\mathrm{dB}}}$ and $dx=\mathrm{ln}\left(10\right)10^{\frac{x+\mu_{\lambda_{\mathrm{dB}}}}{10}}/10~{}d\lambda$, after some manipulation we have
[…]
Appendix C Approximation of $\mu_{\lambda_{\mathrm{dB},i}}$ from the GUE derivative
We approximate the means of individual eigenvalues $\lambda_{\mathrm{dB},i}$ by local maxima of $f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})$. Given that the GUE distribution is positive definite and is zero at the limit $\lambda_{\mathrm{dB}}\rightarrow-\infty$, defining properties of probabilistic distributions, the first $\lambda_{\mathrm{dB},i}$ value for which $f^{\prime}_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})=0$ is a local maximum, and the subsequent root is a local minimum. Maxima and minima alternate in a way such that odd solutions return local maxima. This can be demonstrated by guaranteeing that $f^{\prime\prime}_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})<0$ for the $\lambda_{\mathrm{dB}}$ values that fulfill $f^{\prime}_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})=0$. Expanding on the first derivative from (3), utilizing $\widehat{\lambda}_{\mathrm{dB}}=\lambda_{\mathrm{dB}}-\mu_{\lambda_{\mathrm{dB}}}$, yields
$f^{\prime}_{\widehat{\lambda}_{\mathrm{dB}}}(\widehat{\lambda}_{\mathrm{dB}})=0$ (43)
$\frac{d}{d\widehat{\lambda}_{\mathrm{dB}}}\left[\frac{\alpha_{\lambda_{\mathrm{dB}},D}}{\sigma_{\mathrm{mdg}}}e^{-\frac{(D+1)}{2}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2}}{\sigma_{\mathrm{mdg}}^{2}}}\sum_{k=0}^{D-1}\beta_{\lambda_{\mathrm{dB}},D,k}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2k}}{\sigma_{\mathrm{mdg}}^{2k}}\right]=0$
$\frac{d}{d\widehat{\lambda}_{\mathrm{dB}}}\left[e^{-\frac{(D+1)}{2}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2}}{\sigma_{\mathrm{mdg}}^{2}}}\sum_{k=0}^{D-1}\beta_{\lambda_{\mathrm{dB}},D,k}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2k}}{\sigma_{\mathrm{mdg}}^{2k}}\right]=0.$
Setting $g(\widehat{\lambda}_{\mathrm{dB}})=\sum_{k=0}^{D-1}\beta_{\lambda_{\mathrm{dB}},D,k}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2k}}{\sigma_{\mathrm{mdg}}^{2k}}$ and $\delta_{0}=\frac{D+1}{2\sigma_{mdg}^{2}}$,
$\frac{d}{d\widehat{\lambda}_{\mathrm{dB}}}\left[e^{-\delta_{0}\lambda_{\mathrm{dB}}^{2}}g(\widehat{\lambda}_{\mathrm{dB}})\right]=0$ (44)
$g^{\prime}(\widehat{\lambda}_{\mathrm{dB}})e^{-\delta_{0}\widehat{\lambda}_{\mathrm{dB}}^{2}}-2\delta_{0}\widehat{\lambda}_{\mathrm{dB}}e^{-\delta_{0}\widehat{\lambda}_{\mathrm{dB}}^{2}}g(\widehat{\lambda}_{\mathrm{dB}})=0$
$g^{\prime}(\widehat{\lambda}_{\mathrm{dB}})-2\delta_{0}\widehat{\lambda}_{\mathrm{dB}}g(\widehat{\lambda}_{\mathrm{dB}})=0,$
where $g^{\prime}(\widehat{\lambda}_{\mathrm{dB}})$ is given by
$g^{\prime}(\widehat{\lambda}_{\mathrm{dB}})=\sum_{k=0}^{D-1}\frac{2k\beta_{\lambda_{\mathrm{dB}},D,k}}{\sigma_{\mathrm{mdg}}^{2}}\widehat{\lambda}_{\mathrm{dB}}^{2k-1}.$ (45)
Applying (45) into (44), and replacing $\widehat{\lambda}_{\mathrm{dB}}$ by $\mu_{\lambda_{\mathrm{dB},i}}-\mu_{\lambda_{\mathrm{dB}}}$ yields (5) after some manipulation.
Appendix D Approximation of $\mu_{C_{i}}$ from the derivative of the individual capacities distribution
The distribution for individual capacities is provided in (8). We now assume that a Gaussian distribution can approximate the distribution of individual capacities. If this approximation holds, the mean capacity corresponds to the point where the derivative of the probability density function (PDF) is zero, i.e., the mean can be obtained from the solution to $f^{\prime}_{C_{i}}(c_{i})=0$. To derive this result we consider the following auxiliary elements
$\begin{split}\delta_{1}&=\frac{10\cdot\mathrm{ln}\left(2\right)}{\sigma_{\lambda_{\mathrm{dB},i}}\mathrm{ln}\left(10\right)\sqrt{2\pi}},\\ h_{0}(c_{i})&=\frac{1}{\sigma_{\lambda_{\mathrm{dB},i}}}\left[\frac{10}{\mathrm{ln}\left(10\right)}\mathrm{ln}\left(\frac{2^{c_{i}}-1}{\mathrm{SNR}}\right)-\mu_{\lambda_{\mathrm{dB},i}}\right],\\ h_{1}(c_{i})&=e^{\frac{-h_{0}^{2}(c_{i})}{2}},\\ h_{2}(c_{i})&=2^{c_{i}}h_{1}(c_{i}).\end{split}$ (46)
Applying (7) into (8), and applying the substitutions provided in (46), yields
$f_{C_{i}}(c_{i})=\frac{\delta_{1}}{\left(2^{c_{i}}-1\right)}h_{2}(c_{i}),$ (47)
with the derivative given by
$f^{\prime}_{C_{i}}(c_{i})=\frac{\delta_{1}}{\left(2^{c_{i}}-1\right)^{2}}\left[h^{\prime}_{2}(c_{i})\left(2^{c_{i}}-1\right)-h_{2}(c_{i})\mathrm{ln}\left(2\right)2^{c_{i}}\right].$ (48)
The derivatives of the auxiliary functions in (46) are given by
$\begin{split}h^{\prime}_{0}(c_{i})&=\frac{\delta_{1}\sqrt{2\pi}\cdot 2^{c_{i}}}{\left(2^{c_{i}}-1\right)}\\ h^{\prime}_{1}(c_{i})&=-h_{0}(c_{i})h^{\prime}_{0}(c_{i})e^{\frac{-h_{0}^{2}(c_{i})}{2}}\\ h^{\prime}_{2}(c_{i})&=2^{c_{i}}\left(h^{\prime}_{1}(c_{i})+\mathrm{ln}\left(2\right)h_{1}(c_{i})\right).\end{split}$ (49)
Replacing these terms into (48) results in
$\begin{split}f^{\prime}_{C_{i}}(c_{i})&=\frac{\delta_{1}2^{c_{i}}}{\left(2^{c_{i}}-1\right)^{2}}[\left(h^{\prime}_{1}(c_{i})+\mathrm{ln}\left(2\right)h_{1}(c_{i})\right)\left(2^{c_{i}}-1\right)\\ &~{}~{}-2^{c_{i}}\mathrm{ln}\left(2\right)h_{1}(c_{i})]\\ &=\frac{\delta_{1}2^{c_{i}}}{\left(2^{c_{i}}-1\right)^{2}}\left[h^{\prime}_{1}(c_{i})\left(2^{c_{i}}-1\right)-\mathrm{ln}\left(2\right)h_{1}(c_{i})\right]\\ &=\frac{-\delta_{1}2^{c_{i}}e^{\frac{-h_{0}^{2}(c_{i})}{2}}\mathrm{ln}\left(2\right)}{\left(2^{c_{i}}-1\right)^{2}}\left[\frac{\delta_{1}h_{0}(c_{i})2^{c_{i}}\sqrt{2\pi}}{\mathrm{ln}\left(2\right)}+1\right].\end{split}$ (50)
For $f^{\prime}_{C_{i}}(c_{i})=0$,  (50) can be simplified. Given that the capacity is positive, the first multiplicative terms of (50) are always non-zero, as expressed by
$\frac{-\delta_{1}2^{c_{i}}e^{\frac{-h_{0}^{2}(c_{i})}{2}}\mathrm{ln}\left(2\right)}{\left(2^{c_{i}}-1\right)^{2}}<0,\ \forall c_{i}\in(0,\infty),$ (51)
and therefore, can be disregarded.
After simplifications, replacing the auxiliary terms from (46) into (50), and replacing $c_{i}$ with $\mu_{C_{i}}$, results in (9).
Appendix E Capacity CDF given the semicircular $\lambda_{\mathrm{dB}}$ distribution
The capacity CDF considering the ensemble of eigenvalues can be obtained by integrating (12). Considering $\delta_{2}=\sigma_{\mathrm{mdg}}\mathrm{ln}\left(10\right)/10$, the CDF is given by
[…]
$=$ $\frac{\mathrm{ln}\left(2\right)}{2\pi\delta_{2}}\int_{c_{0}}^{c}\frac{2^{x}}{2^{x}-1}$
$\sqrt{4-\left(\frac{1}{\delta_{2}}\mathrm{ln}\left(\frac{2^{x}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{\sigma_{\mathrm{mdg}}}\right)^{2}}dx,$
where $c_{0}=\mathrm{log}_{2}(\mathrm{SNR}\cdot 10^{\frac{1}{10}\left(\mu_{\lambda_{\mathrm{dB}}}-2\sigma_{\mathrm{mdg}}\right)}+1)$ is the smallest value of $c$ such that $f_{C}(c)$ is strictly real, obtainable from the lowest real-valued boundary of the terms inside the square root. Applying the following variable substitutions
$y$ $=\mathrm{sin^{-1}}\left(\frac{1}{2\delta_{2}}\mathrm{ln}\left(\frac{2^{x}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{2\sigma_{\mathrm{mdg}}}\right)$ (53)
$y_{0}$ $=\mathrm{sin^{-1}}\left(\frac{1}{2\delta_{2}}\mathrm{ln}\left(\frac{2^{c_{0}}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{2\sigma_{\mathrm{mdg}}}\right)=\frac{-\pi}{2}$
$y_{f}$ $=\mathrm{sin^{-1}}\left(\frac{1}{2\delta_{2}}\mathrm{ln}\left(\frac{2^{c}-1}{\mathrm{SNR}}\right)-\frac{\mu_{\lambda_{\mathrm{dB}}}}{2\sigma_{\mathrm{mdg}}}\right)$
$dy$ $=\frac{\mathrm{ln}\left(2\right)\cdot 2^{x}}{\delta_{2}\left(2^{x}-1\right)\sqrt{4-4\cdot\mathrm{sin}^{2}\left(y\right)}}dx,$
results in
[…]
Applying the necessary substitutions from (53) into (54) results in the CDF given in (13).